OpenStudy (anonymous):

Can someone show me how to solve this? If the length of AB is one-fourth the length of AC, what is the length of AB? Length of AC is 30. a)25 b)10 c)30 d)40

OpenStudy (anonymous):

7.5, its not in the choices........ AB= AC/4 and AC =30 so 30/4 =7.5

OpenStudy (anonymous):

so how did you get that? can you explain?

OpenStudy (anonymous):

im a little confused : /

OpenStudy (radar):

In words, AB is 1/4th AC, in algebra AB=AC/4, AC is equal to 30. Substituting AB=30/4 ans is as was stated above 7.5. What do you get when you divide 30 by 4?

OpenStudy (anonymous):

ohh now i understand : ) thanks , you get 7.5